OpenStudy (anonymous):

there are eight contestants in the art contest. How many different ways can first, second, and third place be awarded?

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

@Compassionate @MikeyMaximum @nincompoop

jimthompson5910 (jim_thompson5910):

we have 8 people to pick from for slot 1 how many people do we have to pick from for slot 2?

OpenStudy (anonymous):

i dont know :/

OpenStudy (shamim):

After choosing 1 people frm 8, u hv 7 people to choose for slot 2

jimthompson5910 (jim_thompson5910):

well if one person gets first place, then we have 8-1 = 7 people left, no?

OpenStudy (anonymous):

oh Oops, Yes

jimthompson5910 (jim_thompson5910):

and slot 3?

OpenStudy (anonymous):

6

OpenStudy (anonymous):

So...

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

multiply out those values: 8*7*6 = ??

jimthompson5910 (jim_thompson5910):

the result is the number of ways to pick 3 people from a pool of 8 (order matters)

OpenStudy (anonymous):

336

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (anonymous):

Ok Thx!

jimthompson5910 (jim_thompson5910):

that's the same as the permutation 8 P 3

OpenStudy (shamim):

Why we multiply 8*7*6? Why we dont simply add those 8+7+6?

OpenStudy (shamim):

@AwesomeSays3

OpenStudy (anonymous):

:/ i dont know ask @jim_thompson5910

OpenStudy (shamim):

I know the answer

OpenStudy (anonymous):

what?

OpenStudy (shamim):

I need some time to clear it

OpenStudy (anonymous):

Ok

OpenStudy (shamim):

Just numbering people 1,2,3,4,5,6,7,8

OpenStudy (shamim):

If 1 stood first, second can b 2 or 3 or 4 or 5 or 6 or 7 or 8

OpenStudy (shamim):

So the r 7 ways to choose second if 1 is first. Right?

OpenStudy (shamim):

And if u choose people number 2 for first slot, u will get another 7 ways to choose for slot 2

OpenStudy (shamim):

And if u choose people number 3 for slot 1 , then u will b able to fill up the second slot by another 7 ways

OpenStudy (shamim):

Nd so on

OpenStudy (shamim):

So u will b able to fill up first nd second slot by total number of ways=8*7 Right?

OpenStudy (shamim):

I think u r not following me!

OpenStudy (shamim):

U missed something!!

jimthompson5910 (jim_thompson5910):

No sorry. I don't know what you're asking exactly

OpenStudy (shamim):

I think asker did not understand why u multiply 8*7*6